OpenStudy (anonymous):

simplify using positive exponents p^3q^-1/q^2r^-6

OpenStudy (anonymous):

p^3r^6/q^1+2

OpenStudy (anonymous):

\[\frac{p^3q^{-1}}{q^2r^{-6}}=\frac{p^3r^6}{q^3}\]